# leftmargin=0.15cm, rightmargin=0.15cm, font=italic, vskip=0.10cm, indentfirst=false

### v4

@@ IX Acknowledgments
Technological Factors
Available Sigstore Functionalities 3 – P1, P5, P14
A transparency log, etc P5, P14
Compatibility to other Tools P1
@@
Regulation & Standards 4 – P5, P6, P15, P17
Large User Community 1 – P9
@@
-I Factors Influencing Sigstore Consideration (Non-Sigstore Users).
TABLE XI: Non Sigstore Users: Tooling Issues causing a consideration of Sigstore and other Next-gen tools.
@@
(5) Verification not built in with tool O – P8
@@
(3) Engineers don’t understand the current tool H – P11
@@
-J Additional Discussions & Implications of Results
In addition to the implications discussed in section VI, we outline further implications of our findings below.
-J1 Generalizability of Findings to Other Classes of Tools in Software Engineering
@@
-J2 The Surprising Role of Community in Software Signing Tool Adoption
Many participants cited their involvement in the Sigstore community as a key reason for adopting the tool (Table X). While this may partly reflect our sampling, half of the non-Sigstore users who adopted external tools also emphasized community affiliation as a motivator. Challenges like setting up private Sigstore instances or accessing usage knowledge were frequently linked to the absence of sufficient community support (section V-A2). Other community-related influences, such as the size of the user community and trust in the tool creators, also shaped adoption decisions.
@@
-K Legible Images
We have added a more legible version of our Sigstore workflow imageFigure 3 here Figure 8.

### v5

@@ E.4 Factors Influencing Sigstore Adoption (Sigstore Users)
Steep learning curve P12, P16
Notary Issues 1
Non-demand from customers P12
Technological Factors
Available Sigstore Functionalities 3 – P1, P5, P14
A transparency log, etc P5, P14
Compatibility to other Tools P1
GPG Issues 5
@@
Regulation & Standards (Cross-factor)- M & O 4 – P5, P6, P15, P17
Large Sigstore User Community 1 – P9
@@
E.5 Factors Influencing Sigstore Consideration (Non-Sigstore Users).
Table 12: Non Sigstore Users: Tooling Issues causing a consideration of Sigstore and other Next-gen tools.
@@
(5) Verification not built in with tool T – P8
@@
(3) Engineers don’t understand the current tool P – P11
@@
E.6 Additional Discussions & Implications of Results
In addition to the implications discussed in section˜6, we outline further implications of our findings below.
E.6.1 Generalizability of Findings to Other Classes of Tools in Software Engineering
@@
E.6.2 The Role of Community in Software Signing Tool Adoption
Many participants cited their involvement in the Sigstore community as a key reason for adopting the tool (Table˜11). While this may partly reflect our sampling, half of the non-Sigstore users who adopted external tools also emphasized community affiliation as a motivator. Challenges like setting up private Sigstore instances or accessing usage knowledge were frequently linked to the absence of sufficient community support (section˜5.1.2). Other community-related influences, such as the size of the user community and trust in the tool creators, also shaped adoption decisions.
@@ E.6.2 The Role of Community in Software Signing Tool Adoption
E.6.3 Push-Pull-Barrier Definition
In this appendix, we map our generated themes and subthemes to the Push, Pull, and Barrier factors from the TRL framework by Bansal et al. [10], as described in the results section. The table defines the driver type, corresponding theme, and examples. As this analysis was highly objective, a single rater performed this mapping. The table is here: Table˜13.
Table 13: Driver labels (Push/Pull/Barrier) mapped to concrete codes with factor tags (T/P/O/M) and example participant IDs.
Driver Theme Code Factor Examples
Barrier Policy/ownership Centralize org code signing internally O P8
Barrier Policy/ownership Organization’s open-source policy constraints O P8
Barrier Policy/ownership Organizational use cases don’t fit Sigstore O P8
Barrier Policy/ownership Tool owned by organization O P10
Barrier Risk/privacy/trust Low trust in Sigstore maintainers M P13
Barrier Risk/privacy/trust Sigstore privacy concerns T P13
Barrier Risk/privacy/trust Third-party security risk concerns T P11, P8
Pull Community/engagement Practitioners contribute to Sigstore H P2, P4, P6, P7, P9, P14
Pull Macro pressures Large Sigstore user community M/H P9
Pull Macro pressures Regulation & standards alignment M/O P5, P6, P15, P17
Pull Macro pressures Trust in CNCF/creators M P3
Pull Sigstore capabilities Compatibility with other tools/CI T P1
Pull Sigstore capabilities Transparency log & related features T P5, P14
Push General Key & identity management pain T P5
Push GPG issues (human) Low adoption rates H P1, P14
Push GPG issues (human) Steep learning curve H P12, P16
Push GPG issues (tech) Compatibility with newer technology T P16
Push GPG issues (tech) Key management & other usability pain T P1, P6, P12, P16, P17
Push Notary v1 issues Engineers don’t understand tool P P11
Push Notary v1 issues Key management T P10
Push Notary v1 issues Lack of in-built signing policy T P10
Push Notary v1 issues Lack of regular updates T P12
Push Notary v1 issues Compatibility problems T P12
Push Proprietary/Internal tool issues Difficult to set up T P15
Push Proprietary/Internal tool issues Engineers don’t understand current tool P P11
Push Proprietary/Internal tool issues Integration with other tools/platforms T P8, P11
Push Proprietary/Internal tool issues Non-transparent signing infrastructure T P11
Push Proprietary/Internal tool issues Non-unified implementations across teams O P8, P11
Push Proprietary/Internal tool issues Security concerns about signing infrastructure T P11
Push Proprietary/Internal tool issues Signer ID management problems T P11
Push Proprietary/Internal tool issues Unclear usage info/docs T P11